プリンキピア

第4章離散確率変数

HOMEWORK宿題

Exercises 117–127練習問題 117–127
Exercise 117問題 117
The switchboard in a Minneapolis law office gets an average of 5.5 incoming phone calls during the noon hour on Mondays. Experience shows that the existing staff can handle up to six calls in an hour. Let X = the number of calls received at noon.
英語のヒント
ミネアポリスの法律事務所の電話交換台には、月曜日の正午から1時間に平均5.5件の電話がかかってくる。経験上、現在の職員は1時間に6件まで対応できる。X = 正午からの1時間に受ける電話の件数とする。
  1. Find the mean and standard deviation of X.
    Xの平均と標準偏差を求めよ。
  2. What is the probability that the office receives at most six calls at noon on Monday?
    月曜日の正午からの1時間に受ける電話が6件以下である確率はいくらか。
  3. Find the probability that the law office receives six calls at noon. What does this mean to the law office staff who get, on average, 5.5 incoming phone calls at noon?
    正午からの1時間に電話を6件受ける確率を求めよ。平均5.5件の電話を受ける法律事務所の職員にとって、この結果は何を意味するか。
  4. What is the probability that the office receives more than eight calls at noon?
    正午からの1時間に8件より多くの電話を受ける確率はいくらか。
Solution解き方
  1. X ~ P(5.5); μ = 5.5; σ = 5.5 ≈ 2.3452
  2. P(x ≤ 6) = poissoncdf(5.5, 6) ≈ 0.6860
  3. There is a 15.7% probability that the law staff will receive more calls than they can handle.
    英語のヒント
    事務所の職員が対応できる件数を超える電話を受ける確率は15.7%である。
  4. P(x > 8) = 1 – P(x ≤ 8) = 1 – poissoncdf(5.5, 8) ≈ 1 – 0.8944 = 0.1056
    P(x > 8) = 1 − P(x ≤ 8) = 1 − poissoncdf(5.5, 8) ≈ 1 − 0.8944 = 0.1056
Exercise 118問題 118
The maternity ward at Dr. Jose Fabella Memorial Hospital in Manila in the Philippines is one of the busiest in the world with an average of 60 births per day. Let X = the number of births in an hour.
英語のヒント
フィリピンのマニラにあるホセ・ファベラ記念病院の産科病棟は、1日平均60件の出産があり、世界で最も多忙な産科の一つである。X = 1時間当たりの出生数とする。
  1. Find the mean and standard deviation of X.
    Xの平均と標準偏差を求めよ。
  2. Sketch a graph of the probability distribution of X.
    Xの確率分布のグラフを概略で描け。
  3. What is the probability that the maternity ward will deliver three babies in one hour?
    産科病棟で1時間に3人の赤ちゃんが生まれる確率はいくらか。
  4. What is the probability that the maternity ward will deliver at most three babies in one hour?
    産科病棟で1時間に生まれる赤ちゃんが3人以下である確率はいくらか。
  5. What is the probability that the maternity ward will deliver more than five babies in one hour?
    産科病棟で1時間に5人より多くの赤ちゃんが生まれる確率はいくらか。
Exercise 119問題 119
A manufacturer of Christmas tree light bulbs knows that 3% of its bulbs are defective. Find the probability that a string of 100 lights contains at most four defective bulbs using both the binomial and Poisson distributions.
英語のヒント
クリスマスツリー用電球のメーカーは、製品の3%が不良品だと分かっている。100個の電球をつないだ一連の飾りに、不良電球が4個以下しか含まれない確率を、二項分布とポアソン分布の両方で求めよ。
Solution解き方
Let X = the number of defective bulbs in a string.
X = 一連の飾りに含まれる不良電球の個数とする。
Using the Poisson distribution:
ポアソン分布を使うと、
  • μ = np = 100(0.03) = 3
  • X ~ P(3)
  • P(x ≤ 4) = poissoncdf(3, 4) ≈ 0.8153
Using the binomial distribution:
二項分布を使うと、
  • X ~ B(100, 0.03)
  • P(x ≤ 4) = binomcdf(100, 0.03, 4) ≈ 0.8179
The Poisson approximation is very good—the difference between the probabilities is only 0.0026.
ポアソン近似は非常によく、確率の差はわずか0.0026である。
Exercise 120問題 120
The average number of children a Japanese woman has in her lifetime is 1.37. Suppose that one Japanese woman is randomly chosen.
日本人女性が生涯にもつ子どもの平均人数は1.37人である。日本人女性を1人無作為に選ぶとする。
  1. In words, define the random variable X.
    確率変数Xを言葉で定義せよ。
  2. List the values that X may take on.
    Xがとりうる値を列挙せよ。
  3. Give the distribution of X. X ~ _____(_____,_____)
    Xの分布を示せ。X ~ _____(_____,_____)
  4. Find the probability that she has no children.
    子どもがいない確率を求めよ。
  5. Find the probability that she has fewer children than the Japanese average.
    子どもの人数が日本の平均より少ない確率を求めよ。
  6. Find the probability that she has more children than the Japanese average.
    子どもの人数が日本の平均より多い確率を求めよ。
Exercise 121問題 121
The average number of children a Spanish woman has in her lifetime is 1.47. Suppose that one Spanish woman is randomly chosen.
スペイン人女性が生涯にもつ子どもの平均人数は1.47人である。スペイン人女性を1人無作為に選ぶとする。
  1. In words, define the Random Variable X.
    確率変数Xを言葉で定義せよ。
  2. List the values that X may take on.
    Xがとりうる値を列挙せよ。
  3. Give the distribution of X. X ~ _____(_____,_____)
    Xの分布を示せ。X ~ _____(_____,_____)
  4. Find the probability that she has no children.
    子どもがいない確率を求めよ。
  5. Find the probability that she has fewer children than the Spanish average.
    子どもの人数がスペインの平均より少ない確率を求めよ。
  6. Find the probability that she has more children than the Spanish average .
    子どもの人数がスペインの平均より多い確率を求めよ。
Solution解き方
  1. X = the number of children for a Spanish woman
    X = スペイン人女性の子どもの人数
  2. 0, 1, 2, 3,...
  3. X ~ P(1.47)
  4. 0.2299
  5. 0.5679
  6. 0.4321
Exercise 122問題 122
Fertile, female cats produce an average of three litters per year. Suppose that one fertile, female cat is randomly chosen. In one year, find the probability she produces:
英語のヒント
繁殖可能な雌猫は、1年に平均3回出産する。そのような雌猫を1匹無作為に選ぶとする。1年間について、次の確率を求めよ。
  1. In words, define the random variable X.
    確率変数Xを言葉で定義せよ。
  2. List the values that X may take on.
    Xがとりうる値を列挙せよ。
  3. Give the distribution of X. X ~ _______
    Xの分布を示せ。X ~ _______
  4. Find the probability that she has no litters in one year.
    1年間に一度も出産しない確率を求めよ。
  5. Find the probability that she has at least two litters in one year.
    1年間に少なくとも2回出産する確率を求めよ。
  6. Find the probability that she has exactly three litters in one year.
    1年間にちょうど3回出産する確率を求めよ。
Exercise 123問題 123
The chance of having an extra fortune in a fortune cookie is about 3%. Given a bag of 144 fortune cookies, we are interested in the number of cookies with an extra fortune. Two distributions may be used to solve this problem, but only use one distribution to solve the problem.
英語のヒント
フォーチュンクッキーに、おみくじの紙が余分に1枚入っている確率は約3%である。144個入りの袋で、余分な紙が入ったクッキーの個数に関心がある。この問題には二つの分布を使えるが、どちらか一つだけを使って解け。
  1. In words, define the random variable X.
    確率変数Xを言葉で定義せよ。
  2. List the values that X may take on.
    Xがとりうる値を列挙せよ。
  3. Give the distribution of X. X ~ _____(_____,_____)
    Xの分布を示せ。X ~ _____(_____,_____)
  4. How many cookies do we expect to have an extra fortune?
    余分なおみくじが入ったクッキーは何個と予想されるか。
  5. Find the probability that none of the cookies have an extra fortune.
    余分なおみくじが入ったクッキーが一つもない確率を求めよ。
  6. Find the probability that more than three have an extra fortune.
    余分なおみくじが入ったクッキーが3個より多い確率を求めよ。
  7. As n increases, what happens involving the probabilities using the two distributions? Explain in complete sentences.
    nが増えると、二つの分布で求めた確率はどうなるか。完全な文で説明せよ。
Solution解き方
  1. X = the number of fortune cookies that have an extra fortune
    X = 余分なおみくじが入ったフォーチュンクッキーの個数
  2. 0, 1, 2, 3,... 144
  3. X ~ B(144, 0.03) or P(4.32)
    X ~ B(144, 0.03)またはP(4.32)
  4. 4.32
  5. 0.0124 or 0.0133
    0.0124または0.0133
  6. 0.6300 or 0.6264
    0.6300または0.6264
  7. As n gets larger, the probabilities get closer together.
    nが大きくなるにつれ、確率は互いに近づく。
Exercise 124問題 124
For every 200 U.S. kids and teens, the average number who suffer from obsessive compulsive disorder is one. Out of a randomly chosen group of 600 U.S. kids and teens determine the following.
英語のヒント
米国の子どもや10代の若者200人当たり、強迫性障害に苦しむ人は平均1人である。無作為に選んだ600人について、次を求めよ。
  1. In words, define the random variable X.
    確率変数Xを言葉で定義せよ。
  2. List the values that X may take on.
    Xがとりうる値を列挙せよ。
  3. Give the distribution ofX. X ~ _____(_____,_____)
    Xの分布を示せ。X ~ _____(_____,_____)
  4. How many are expected to suffer from obsessive compulsive disorder?
    強迫性障害に苦しむ人は何人と予想されるか。
  5. Find the probability that no one suffers from obsessive compulsive disorder.
    強迫性障害に苦しむ人が1人もいない確率を求めよ。
  6. Find the probability that more than four suffer from obsessive compulsive disorder.
    強迫性障害に苦しむ人が4人より多い確率を求めよ。
Exercise 125問題 125
The chance of an IRS audit for a tax return with over $25,000 in income is about 2% per year. Suppose that 100 people with tax returns over $25,000 are randomly picked. We are interested in the number of people audited in one year. Use a Poisson distribution to anwer the following questions.
所得が25,000ドルを超える申告書がIRSの税務調査を受ける確率は、年間約2%である。そのような申告者100人を無作為に選ぶとする。1年間に調査を受ける人数に関心がある。ポアソン分布を使って次の問いに答えよ。
  1. In words, define the random variable X.
    確率変数Xを言葉で定義せよ。
  2. List the values that X may take on.
    Xがとりうる値を列挙せよ。
  3. Give the distribution of X. X ~ _____(_____,_____)
    Xの分布を示せ。X ~ _____(_____,_____)
  4. How many are expected to be audited?
    何人が税務調査を受けると予想されるか。
  5. Find the probability that no one was audited.
    税務調査を受ける人が1人もいない確率を求めよ。
  6. Find the probability that at least three were audited.
    少なくとも3人が税務調査を受ける確率を求めよ。
Solution解き方
  1. X = the number of people audited in one year
    X = 1年間に税務調査を受ける人数
  2. 0, 1, 2, ..., 100
  3. X ~ P(2)
  4. 2
  5. 0.1353
  6. 0.3233
Exercise 126問題 126
Approximately 8% of students at a local high school participate in after-school sports all four years of high school. A group of 60 seniors is randomly chosen. Of interest is the number that participated in after-school sports all four years of high school.
ある地元の高校では、生徒の約8%が高校の4年間すべてで放課後のスポーツ活動に参加する。最上級生60人を無作為に選ぶ。4年間すべて参加した人数に関心がある。
  1. In words, define the random variable X.
    確率変数Xを言葉で定義せよ。
  2. List the values that X may take on.
    Xがとりうる値を列挙せよ。
  3. Give the distribution of X. X ~ _____(_____,_____)
    Xの分布を示せ。X ~ _____(_____,_____)
  4. How many seniors are expected to have participated in after-school sports all four years of high school?
    4年間すべてで放課後のスポーツ活動に参加した最上級生は何人と予想されるか。
  5. Based on numerical values, would you be surprised if none of the seniors participated in after-school sports all four years of high school? Justify your answer numerically.
    数値に基づくと、4年間すべてで放課後のスポーツ活動に参加した最上級生が1人もいなかったら驚くか。数値で根拠を示せ。
  6. Based on numerical values, is it more likely that four or that five of the seniors participated in after-school sports all four years of high school? Justify your answer numerically.
    数値に基づくと、4年間すべてで放課後のスポーツ活動に参加した最上級生が4人の場合と5人の場合では、どちらの可能性が高いか。数値で根拠を示せ。
Exercise 127問題 127
On average, Pierre, an amateur chef, drops three pieces of egg shell into every two cake batters he makes. Suppose that you buy one of his cakes.
英語のヒント
アマチュア料理人のピエールは、平均してケーキの生地2個分を作るごとに、卵の殻を3片落としてしまう。彼のケーキを1個買うとする。
  1. In words, define the random variable X.
    確率変数Xを言葉で定義せよ。
  2. List the values that X may take on.
    Xがとりうる値を列挙せよ。
  3. Give the distribution of X. X ~ _____(_____,_____)
    Xの分布を示せ。X ~ _____(_____,_____)
  4. On average, how many pieces of egg shell do you expect to be in the cake?
    ケーキには平均して何片の卵の殻が入っていると予想されるか。
  5. What is the probability that there will not be any pieces of egg shell in the cake?
    ケーキに卵の殻が全く入っていない確率はいくらか。
  6. Let’s say that you buy one of Pierre’s cakes each week for six weeks. What is the probability that there will not be any egg shell in any of the cakes?
    ピエールのケーキを毎週1個ずつ6週間買うとする。どのケーキにも卵の殻が入っていない確率はいくらか。
  7. Based upon the average given for Pierre, is it possible for there to be seven pieces of shell in the cake? Why?
    英語のヒント
    ピエールについて与えられた平均に基づくと、ケーキに殻が7片入っていることはありうるか。理由も述べよ。
Solution解き方
  1. X = the number of shell pieces in one cake
    X = ケーキ1個に含まれる殻のかけらの数
  2. 0, 1, 2, 3,...
  3. X ~ P(1.5)
  4. 1.5
  5. 0.2231
  6. 0.0001
  7. Yes
    はい。